# TrAME: Trajectory-Anchored Multi-View Editing for Text-Guided 3D Gaussian Splatting Manipulation

### v1

@@ 4.3 View-Consistent Attention Control
We utilize a source branch and a target branch to achieve 3D consistent editing. The source branch takes in unedited original views, acting as a reference for the target branch, providing structural and semantic guidance for generating edited views. Previous studies have noted that the structural layout of an image is formed in the early stages of diffusion and is closely linked to self-attention queries. To ensure that the edited views produced by the target branch maintain structural multi-view consistency, we opt to inject self-attention queries from the source branch into the target branch during the early stages of diffusion (mid-bottom part of Figure 1). Specifically, denote the attention calculation as:
@@
Maintaining structural multi-view consistency alone is insufficient; semantic multi-view consistency is also crucial. To achieve this, we implemented KV propagation and KV reference mechanisms (bottom-left of Figure 1). Given a sequence of $N$ views $\{{\bm{I}}^{n}\}_{n=1}^{N}$ resulting from consecutive camera motion, we divide the sequence into equal-partitioned context segments $\{{\bm{C}}^{i}\}_{i=1}^{L}$. Each context ${\bm{C}}^{i}$ has a length of $N/L$ frames and contains a keyframe ${\bm{I}}_{1}^{i}$. Views within the same context exhibit minimal movement or view shift, their semantics are relatively consistent. For views within the same context, we utilized an efficient KV propagation method, where the self-attention keys and values of the keyframe $\{{\bm{K}}_{1}^{i},{\bm{V}}_{1}^{i}\}$ of $i$-th context is directly propagated to the rest of the frames in the same context. Denote the output of self-attention of $j$-th frame in context $i$ as ${\bm{O}}^{i}_{j}$, we can formulate this attention propagation as:
@@
Significant view shifts can occur across different contexts, leading to noticeable semantic differences. Applying KV propagation on keyframes across different contexts would accumulate significant distortion. To mitigate this, we employ KV reference for keyframes across contexts. This involves inflating the self-attention keys and values across keyframes of different contexts and processing them jointly. The self-attention keys and values from other keyframes serve as a mutual reference for the semantic information of the edited entity. For self-attention keys $\{{\bm{K}}^{i}\}_{i=1}^{l}$ and values $\{{\bm{V}}^{i}\}_{i=1}^{l}$ from $l$ different keyframes, we concatenate them respectively to form cross-context KV pairs. The operation is formulated as:
@@
To maintain fidelity with the original images in the non-edited areas, we applied a cross-attention control (bottom-right of Figure 1) and local blending strategy similar to Prompt-to-Prompt [14] and InfEdit [34]. Cross-attention control is done by replacing the attention maps corresponding to non-edited tokens in the target branch with those from the source branch. We use the semantic mask ${\bm{M}}$ obtained from semantic tracing [5] to blend noisy target latent with source latent, formulated as:
@@
Table 1: Quantitative Results. Compared to DDS and GaussianEditor, our method outperforms across evaluated metrics.
Methods Reference CLIP Score [2] CLIP Directional Score [2] User Study
DDS [13] ICCV (2023) 0.2242 0.1336 5.59%
GaussianEditor [5] CVPR (2024) 0.2343 0.1548 25.47%
TrAME (Ours) Ours 0.2550 0.2002 68.94%
Figure 3: Qualitative comparison. A comparative analysis of experimental results for single-scene editing across State-of-the-Art methods and ours under identical conditions.
5 Experiments
In this section, we provide a comprehensive overview of our experimental setup, which encompasses the testing data, baseline methods and implementation details. We examine the effectiveness of our method with quantitative and qualitative results. Additionally, we conduct ablation studies on each proposed component to dissect their individual contributions to the overall performance.
5.1 Implementation Details
@@
5.2 Baseline Methods
We compare our method with two state-of-the-art 3DGS editing methods: (1) Delta Distillation Sampling (DDS)[13], which represents the SDS optimization strategy for editing and uses an implicit score function loss derived from a diffusion model as guidance; (2) GaussianEditor [5], which represents an editing method based on Instruct-P2P [2], guiding the editing of the 3D Gaussian Splatting (3DGS) by iteratively updating the training set. Note that other 3D scene editing methods either lack publicly available code or utilize 3D representations other than 3DGS. We conduct a quantitative comparison using the CLIP Score and CLIP Directional Score [2] to evaluate the visual resemblance between the edited 3DGS and the given prompt. Additionally, we carry out a user study aimed at assessing the effectiveness of these editing methods from a human perspective.
5.3 Quantitative Results
@@
In our quantitative analysis, we employ the CLIP Score and CLIP Directional Score [2] to assess the alignment of the edited 3D models with the target text prompts on the rendered views of the edited 3DGS. These results indicate that our 3D editing method is effective at consolidating user edits into 3D scenes. To further validate our approach, we organized a user study involving 23 participants over 14 different cases, where users were asked to evaluate and choose the best edits based on image quality, alignment with the target prompt, and consistency across different views. The quantitative evaluation results, presented in Table 1, compare our method against baseline methods across various test scenes. Our method excels in CLIP Score, CLIP Directional Score, and user study scores, confirming its superior editing outcomes.
5.4 Qualitative Results
@@
As shown in Figure 3 and Figure 4, our qualitative analysis showcases that our method ensures multi-view consistency and fine-grained details in 3D scene editing. We display outcomes from various perspectives in different scenes, underscoring the effectiveness of our method. Based on the results, we can observe that DDS faces challenges in achieving detailed and faithful editing. GaussianEditor struggles with maintaining multi-view consistency, as seen in inconsistencies such as the hair color of the elf and the nose of the clown when viewed from different angles, and in consolidating fine-grained details onto 3D textures, such as accumulated snow on park benches and patterns on marble tables (examples shown in the second and third columns of Figure 4). This shortcoming of the compared methods becomes particularly apparent when trying to capture the intricate patterns and fine details inherent to such materials, highlighting a critical area where our approach excels in producing view-consistent and detailed 3D scene edits.
@@
Figure 4: Qualitative comparison. A comparative analysis of experimental results for various scenes encompassing both partial and global editing under identical conditions.
5.5 Ablation Studies
Study on VCAC. The results presented in Figure 5 demonstrate that the VCAC module effectively enhances cross-view structural and semantic consistency. The views modified by our VCAC module align more accurately with the original views in terms of structural elements like poses and facing directions, and they also demonstrate better semantic coherence, exhibiting similar appearances. In particular, the outputs generated by our VCAC module maintain a uniform appearance across various views, with the facing direction of Spiderman in the figures aligning cohesively with the pose of the man in the original views.
Study on TAS. The ablated results obtained without TAS were derived by applying 3D editing solely based on the final obtained edited views, without feedback from the 3D Gaussians. The results depicted in Figure 6 show that TAS produces more naturalistic and consistent appearances, while the ablated results exhibit undesirable blurry artifacts and inconsistencies across views. Specifically, the hair color around the clown’s forehead in the ablated results varies across views, and the lips of the clown show ghosting shadows.
@@ 5.5 Ablation Studies
Figure 5: Ablation study on the VCAC module for 2D view editing. The results display edited views with (w. VCAC) and without the VCAC module (w/o VCAC).
@@
6 Conclusion and Future Work
Our research introduces TrAME, a progressive 3DGS editing framework that effectively improves multi-view consistency via Trajectory Anchored Scheme (TAS) and View-Consistent Attention Control (VCAC) module. Our work bridges the gap between optimization-based and reconstruction-based editing methods, offering a unified perspective for selecting better design choices in 3D scene editing methods. Our limitations reside in complex view-consistent edits with large object deformations, which will be the focus of our future work.

### v2

@@ IV-C View-Consistent Attention Control
We utilize a source branch and a target branch to achieve 3D consistent editing. The source branch takes in unedited original views, acting as a reference for the target branch, providing structural and semantic guidance for generating edited views. Previous studies have noted that the structural layout of an image is formed in the early stages of diffusion and is closely linked to self-attention queries. To ensure that the edited views produced by the target branch maintain structural multi-view consistency, we opt to inject self-attention queries from the source branch into the target branch during the early stages of diffusion (mid-bottom part of Fig. 1). Specifically, denote the attention calculation as:
@@
Maintaining structural multi-view consistency alone is insufficient; semantic multi-view consistency is also crucial. To achieve this, we implemented KV propagation and KV reference mechanisms (bottom-left of Fig. 1). Given a sequence of $N$ views $\{{\bm{I}}^{n}\}_{n=1}^{N}$ resulting from consecutive camera motion, we divide the sequence into equal-partitioned context segments $\{{\bm{C}}^{i}\}_{i=1}^{L}$. Each context ${\bm{C}}^{i}$ has a length of $N/L$ frames and contains a keyframe ${\bm{I}}_{1}^{i}$. Views within the same context exhibit minimal movement or view shift, their semantics are relatively consistent. For views within the same context, we utilized an efficient KV propagation method, where the self-attention keys and values of the keyframe $\{{\bm{K}}_{1}^{i},{\bm{V}}_{1}^{i}\}$ of $i$-th context is directly propagated to the rest of the frames in the same context. Denote the output of self-attention of $j$-th frame in context $i$ as ${\bm{O}}^{i}_{j}$, we formulate this attention propagation as:
@@
Significant view shifts occur across different contexts, leading to noticeable semantic differences. Applying KV propagation on keyframes across different contexts would accumulate significant distortion. To mitigate this, we employ KV reference for keyframes across contexts. This involves inflating the self-attention keys and values across keyframes of different contexts and processing them jointly. The self-attention keys and values from other keyframes serve as a mutual reference for the semantic information of the edited entity. For self-attention keys $\{{\bm{K}}^{i}\}_{i=1}^{l}$ and values $\{{\bm{V}}^{i}\}_{i=1}^{l}$ from $l$ different keyframes, we concatenate them respectively to form cross-context KV pairs. The operation is formulated as:
@@
To maintain fidelity with the original images in the non-edited areas, we apply a cross-attention control (bottom-right of Fig. 1) and local blending strategy similar to Prompt-to-Prompt [50] and InfEdit [35]. Cross-attention is done by replacing the attention maps corresponding to non-edited tokens in the target branch with those from the source branch. We use the semantic mask ${\bm{M}}$ obtained from semantic tracing [34] to blend noisy target latent with source latent, formulated as:
@@
V Experiments
TABLE II: Quantitative Results. Comparative analysis on the reconstruction-based methods regarding the coherence between 2D view edits and 3D scene modifications.
Methods PSNR $\uparrow$ SRE $\uparrow$ RMSE (×$\times 10^{-5}$) $\downarrow$ CLIP Image-to-Image Similarity $\uparrow$
ViT-B/16 ViT-B/32 ViT-L/14
IN2N [49] 87.860 $\pm$1.050 29.828 $\pm$0.771 4.438 $\pm$0.509 0.888 $\pm$0.036 0.896 $\pm$0.029 0.866 $\pm$0.033
VICA [41] 88.846 $\pm$0.676 30.369 $\pm$0.380 3.892 $\pm$0.246 0.880 $\pm$0.024 0.867 $\pm$0.023 0.858 $\pm$0.021
GSEditor [34] 96.040 $\pm$1.755 34.487 $\pm$0.991 1.649 $\pm$0.357 0.939 $\pm$0.016 0.936 $\pm$0.017 0.914 $\pm$0.026
TrAME (w/o TAS & w/o VCAC) 87.214 $\pm$1.444 30.575 $\pm$0.713 4.438 $\pm$0.720 0.669 $\pm$0.053 0.675 $\pm$0.066 0.601 $\pm$0.053
TrAME (w/o TAS & w. VCAC) 90.101 $\pm$0.988 32.283 $\pm$0.581 3.169 $\pm$0.350 0.801 $\pm$0.061 0.804 $\pm$0.058 0.754 $\pm$0.061
TrAME (w. TAS & w. VCAC) 102.449 $\pm$0.499 38.322 $\pm$0.247 0.758 $\pm$0.044 0.962 $\pm$0.005 0.978 $\pm$0.003 0.937$\pm$0.011
In this section, we provide a comprehensive overview of our experimental setup, including the testing data, baseline methods, and implementation details. We evaluate the effectiveness of our method using both quantitative and qualitative results. Additionally, we conduct ablation studies on each proposed component to analyze their individual contributions to overall performance. In the final part, we further analyze the limitations and broader impact of our work.
V-A Implementation Details
@@
V-B Baseline Methods
We conduct a comparative analysis of our method against five existing 3D scene editing approaches. Specifically, Instruct-NeRF2NeRF (IN2N) [49], ViCA-NeRF (VICA) [41], and Posterior Distillation Sampling (PDS) [26] employ NeRF as a 3D representation, while GaussianEditor (GSEditor) [34] and Delta Denoising Score (DDS) [31] utilize 3DGS for this purpose. Furthermore, PDS and DDS serve as optimization-based 3D editing baselines, whereas the remaining methods are reconstruction-based. Our quantitative evaluation is based on the CLIP Score and CLIP Directional Score [15], assessing the semantic coherence between the edited 3DGS and the given prompt. In addition, for the reconstruction-based methods, we measure image-to-image similarity between the edited and rendered views of the scenes, to evaluate the consistency between 2D view edits and 3D scene modifications.
TABLE III: Quantitative Results: CLIP Directional Score against NeRF-based methods (top) and 3DGS-based methods (middle).
Methods CLIP Directional Score
ViT-B/16 $\uparrow$ ViT-B/32 $\uparrow$ ViT-L/14 $\uparrow$
IN2N [49] 0.148 $\pm$0.068 0.149 $\pm$0.060 0.127 $\pm$0.064
VICA [41] 0.157 $\pm$0.052 0.151 $\pm$0.044 0.145 $\pm$0.044
PDS [26] 0.163 $\pm$0.080 0.160 $\pm$0.075 0.155 $\pm$0.066
DDS [31] 0.121 $\pm$0.074 0.134 $\pm$0.075 0.111 $\pm$0.071
GSEditor [34] 0.166 $\pm$0.056 0.162 $\pm$0.055 0.161 $\pm$0.054
TrAME (Ours) 0.199 $\pm$0.054 0.201 $\pm$0.055 0.186 $\pm$0.049
Figure 3: Qualitative comparison. A comparative analysis of experimental results for single-scene editing across State-of-the-Art methods and ours. Please zoom in for more geometry and textural details.
Figure 4: Qualitative comparison. A comparative analysis of different design choices of DDCM coefficient $\kappa$ on 2D view editing.
Figure 5: Qualitative results under different hyper-parameter settings. Rendered RGB and depth representations illustrating the implications of different DDCM coefficient $\kappa$ on 3DGS editing.
V-C Quantitative Results
@@
In our quantitative analysis, we employ the CLIP Score and CLIP Directional Score [15] to assess the alignment of the edited 3D models with the target text prompts on the rendered views of the edited 3DGS. These results indicate that our 3D editing method is effective at consolidating user edits into 3D scenes. The quantitative evaluation results, presented in Table III, compare our method against baseline methods across various test scenes. Our method excels in CLIP Score, CLIP Directional Score, confirming its superior editing outcomes.
Additionally, for reconstruction-based methods, we assess the coherence between 2D view edits and the rendered views of the modified 3D scenes by calculating image-to-image similarity. This is achieved by measuring peak signal-to-noise ratio (PSNR), signal-to-reconstruction error ratio (SRE), root mean square error (RMSE), and CLIP image-to-image similarity using three different backbones: ViT-B/16, ViT-B/32, and ViT-L/14. The results depicted in Table II demonstrate that our method achieves greater consistency between 2D view edits and 3D scene modifications, underscoring the effectiveness of our approach.
V-D Qualitative Results
@@
As shown in Fig. 3 and Fig. 6, our qualitative analysis showcases that our method ensures multi-view consistency and fine-grained details in 3D scene editing. We display outcomes from various perspectives in different scenes, underscoring the effectiveness of our method. The qualitative results reveal that IN2N, VICA and PDS struggle to produce accurate and localized edits, while DDS faces challenges in achieving detailed and faithful editing. VICA, in particular, often generates blurry results due to the errors of depth estimation. PDS occasionally fails to produce correct scene colors, as depicted in the “Hulk” case, “Golden bear” case and the “Park” case in Fig. 6. Moreover, the inconsistent view edits of GSEditor often result in unstable optimization, leading to visible artifacts, such as the noisy colorization of the man’s jacket in Fig. 3 and the “autumn park” case in Fig. 6. These shortcomings of the compared methods become particularly apparent when trying to capture the intricate patterns and fine details inherent to such materials, highlighting a critical area where our method excels in producing view-consistent and detailed 3D scene edits.
@@
Figure 6: Qualitative comparison. A comparative analysis of experimental results for various scenes encompassing both partial and global editing. Please zoom in for more geometry and textural details.
V-E Ablation Studies
@@
Figure 7: Ablation study on the VCAC module for 2D view editing. The results display edited views with (w.) and without the VCAC module (w/o).
@@
Study on VCAC. The qualitative results presented in Fig. 7 demonstrate that the VCAC module effectively enhances cross-view structural and semantic consistency. The views modified by our VCAC module align more accurately with the original views in terms of structural elements like poses and facing directions, and they also demonstrate better semantic coherence, exhibiting similar appearances. In particular, the outputs generated by our VCAC module maintain a uniform appearance across various views, with the facing direction of Spiderman in the figures aligning cohesively with the pose of the man in the original views. Additionally, the quantitative results in Table II indicate that the VCAC module significantly reduces inconsistencies between 2D edited views and 3D scene modifications by producing consistent 2D view edits, thereby further validating its effectiveness in preserving coherence across different views.
Study on TAS. The ablated results obtained without TAS were derived by applying 3D editing solely based on the final obtained edited views, without feedback from the 3D Gaussians. The qualitative results depicted in Fig. 8 show that TAS produces more naturalistic and consistent appearances, while the ablated results exhibit undesirable blurry artifacts and inconsistencies across views. Specifically, the hair color around the clown’s forehead in the ablated results varies across views, and the lips of the clown show ghosting shadows. Moreover, the quantitative results in Table II confirm that our proposed TAS effectively mitigates accumulated errors during the 2D view editing process, thereby resulting in higher consistency with the rendered views of the modified 3D scenes.
Study on DDCM coefficient $\kappa$. We conducted a qualitative analysis of the effects of varying $\kappa$ values on both 2D view and 3D scene editing. The results are presented in Fig. 4 and Fig. 5. As shown in Fig. 4, selecting a $\kappa$ value greater or equal to 1.0 often results in over-saturated colors and excessively sharpened edges, thereby reducing the fidelity of the edited image. The optimal value of $\kappa$ was determined to be around 0.95, which yields the highest visual quality. Additionally, we observed that using a larger $\kappa$ in later stages of editing may hinder the ability to properly edit geometry. For instance, in Fig. 5, the elf’s right ear does not consolidate effectively with $\kappa=0.9$. Hence, a smaller $\kappa$ should be applied during the later stages of editing.
Study on query injection threshold $t_{q}$. We quantitatively analyze the effect of varying self-attention query injection thresholds. The results, presented in Table IV, reveal that performance is optimal around $t_{q}=0.5$ and deteriorates as $t_{q}$ increases. This decline is attributed to the over-constraint imposed by query injection. Specifically, query injection tends to impose excessive constraints on the edited images, resulting in reduced modifications to the edited views, which eventually leads to diminished overall editing effectiveness.
V-F Limitations and Broader impact
Limitations. Our limitations lie in performing complex, view-consistent edits involving significant object deformations. As illustrated in Fig. 9, both our approach and existing baselines encounter difficulties with scene edits that necessitate extensive deformation or displacement. This limitation is primarily due to the lack of 3D priors in 2D image editing models and the insufficient integration of 3D geometry required to achieve the desired edits. In future work, we aim to leverage the explicit geometry of the 3D Gaussian Splatting to incorporate multi-level hierarchical editing guidance, facilitating the execution of complex edits in a coarse-to-fine manner.
Broader impact. Our work has the potential to significantly advance the creation of immersive virtual environments, thereby enhancing applications in gaming, virtual reality, and film production. However, the capability to seamlessly edit and manipulate 3D scenes also carries the risk of being exploited to create misleading or harmful visual content. This is particularly concerning in contexts where authenticity is crucial, such as in news media or scientific visualization.
TABLE IV: Quantitative experiment for evaluating the CLIP Directional Score on rendered views of edited 3D scenes with varying values of $t_{q}$.
CLIP Directional Score
ViT-B/16 $\uparrow$ ViT-B/32 $\uparrow$ ViT-L/14 $\uparrow$
$t_{q}=0.4$ 0.222$\pm$0.049 0.191$\pm$0.065 0.195$\pm$0.053
$t_{q}=0.5$ 0.224$\pm$0.051 0.188$\pm$0.067 0.195$\pm$0.052
$t_{q}=0.6$ 0.215$\pm$0.054 0.182$\pm$0.072 0.185$\pm$0.059
$t_{q}=0.7$ 0.203$\pm$0.063 0.180$\pm$0.080 0.171$\pm$0.078
$t_{q}=0.8$ 0.191$\pm$0.069 0.166$\pm$0.083 0.162$\pm$0.081
$t_{q}=0.9$ 0.173$\pm$0.084 0.149$\pm$0.089 0.145$\pm$0.086
Figure 9: Failure cases. Our methods and existing baselines fails to perform complex edits that require large object deformations.
VI Conclusion
Our research introduces TrAME, a progressive 3DGS editing framework that effectively improves multi-view consistency via Trajectory Anchored Scheme (TAS) and View-Consistent Attention Control (VCAC) module. Our work bridges the gap between optimization-based and reconstruction-based editing methods, offering a unified perspective for selecting better design choices in 3D scene editing methods.